# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ 4.6. Cross-domain Recommendation
Cross-domain recommendation (CDR) is one of the representative methods to alleviate the data sparsity problem and the cold-start problems in traditional RS with the auxiliary user behaviors collected from other domains (Cao et al., 2022; Ma et al., 2024b). It aims to leverage the correlation of users’ multi-domain behaviors to accurately transfer the informative knowledge from the source domains and model users’ personalized preference in the target domain (Gao et al., 2021a; Bi et al., 2020). Traditional CDR methods typically conduct the cross-domain knowledge transfer by modeling the feature-level cross-domain correlations with aligned constraint (Krishnan et al., 2020; Liu et al., 2020), adversarial training (Liu et al., 2021a) and contrastive learning (Zheng et al., 2022; Ma et al., 2024b). However, these methods merely focus on the randomly-selected negative samples from the target domain (Krishnan et al., 2020; Liu et al., 2020; Ma et al., 2024b; Liu et al., 2021a), ignoring the cross-domain differences at the sample level. This may disregard users’ preference discrepancy between multiple domains and leads to the sub-optimal performance. DDGHM (Zheng et al., 2022) utilizes random mask operator for negative pair augmentation in CL optimization. Nevertheless, when it comes to recommendation tasks, the approach of random sampling within the target domain is still employed. RealHNS (Ma et al., 2023) is the pioneer to systematically discover the false and refine the real from all HNS in CDR. It conducts the multi-grained HNS selectors to find high-quality general and cross-domain HNS with a curriculum learning framework to alleviate cross-domain negative transfer. Notably, there exists substantial uncharted territories for negative sampling methods in CDR, which awaits the investigation by future researchers. This encompasses various areas, including, but not limited to, KNS, ANG, and Debiased IRW.
@@ 4.7. CL-enhanced Recommendation
Inspired by the success of Contrastive Learning (CL) in Computer Vision (Han et al., 2021; Dai and Lin, 2017), Natural Language Processing (Kim et al., 2021b; Hassani and Khasahmadi, 2020) and Graph Representation Learning (Mo et al., 2022; Gao et al., 2021b), some works (Shuai et al., 2022; Wang et al., 2022; Zhou et al., 2020) apply CL paradigm in RS. CL aims to maximize the agreement between positive pairs (an item’s representation and its corresponding augmented representations) and minimize it between negative pairs (an item’s representation and other items’ augmented representations) to acquire the transferable knowledge between the original view and the augmented view. Intuitively, CL relies heavily on its negative augmentation (NA) strategy, which is due to the fact that the appropriate augmented view can facilitate the recommender to exploit richer underlying semantic correlation. The most widely used augmentation strategy in CL-based recommendation is the in-batch negative sampling. It treats the augmented representations of the remaining items in the mini-batch where the positive example is located as negative examples, rather than selecting from the corpus, thereby aiding in accelerating its training process without the imposition of additional computational complexity.
@@
To deeply understand the negative augmentation strategies in CL-based recommendation, we comprehensively dissect its core methodology into four distinct dimensions: pre-defined NA, graph-based NA, sequence-based NA and scenarios-aware NA. Firstly, pre-defined NA fundamentally aims to refine the CL process by augmenting the negative sample pool based on the pre-defined criteria (Du et al., 2022; Zhang et al., 2021b; Gu et al., 2022), which are typically derived from domain-specific expertise or task-specific benchmarks. For instance, CauseRec (Zhang et al., 2021b) models the counterfactual data distribution by replacing the identified indispensable concepts in user behavioral sequence as the negative pairs. Secondly, graph-based NA refers to generating the negative pairs based on distinct types of graph structures, including interaction graphs, social graphs, and heterogeneous graphs (Zhang et al., 2021a; Wu et al., 2021a; Yang et al., 2021d; Ma et al., 2022). It typically perturbs the nodes, edges, and graph structures to provide a nuanced understanding of user-item interactions. SGL (Wu et al., 2021a), EGLN (Yang et al., 2021d), and RGCL (Shuai et al., 2022) fully consider the correlations between nodes and edges within the graph structure to propose several graph-based NA strategies, including node (edge) dropout, node (edge) discrimination, random walk, and structure perturbation. Thirdly, sequence-based NA focuses on enhancing the CL optimization by augmenting negative samples (e.g., shuffle, mask, crop, etc.) derived from historical behavior sequences (Zhou et al., 2020; Tian et al., 2022; Xie et al., 2022a; Wang et al., 2022). Its effective mechanism resides in comprehensively capturing the inherent dynamic characteristics of user behavior sequences. With the tailored self-supervised representation augmentation, it effectively handles the sequence structure and builds more robust and accurate recommenders. Finally, scenarios-aware NA involves tailoring the augmentation of negative samples to accommodate the intricacies and demands of various recommendation scenarios. Here, MMR derives multi-view relationships of the user’s preference and item’s representation across multiple modalities with CL to avoid the problem of limited diversity or bias (Wei et al., 2021; Tao et al., 2022), and MBR constructs negative examples based on the auxiliary behavior data to model the different user behavior patterns through CL (Gu et al., 2021; Yang et al., 2021a). Adaptively selecting the appropriate NA strategy based on the scenario requirements and data availability in CL-based recommendation tasks is sensible. Meanwhile, how to construct a universal augmentation method to serve various recommendation scenarios remains an urgent issue.
@@ 5.3. Causal Inference for Negative Sample Understanding
Causality is the science of cause and effect in which the effect partly originates from the cause to some extent. As mentioned in (Gao et al., 2022b), causal inference refers to the process of determining and further leveraging the causal relation from the experimental results or observational data. In the recent past, causal inference has been introduced into RS (Yang et al., 2021b) to extract the causality within the recommenders in modeling intricate user preferences, departing from the traditional correlation-centric paradigm (He et al., 2022, 2023b). This paradigm aims to determine how user historical behaviors influence the recommendation results, allowing for a comprehensive comprehension of user preferences.
From the perspective of negative sampling, we posit that the integration of negative sampling with causal inference holds the potential to enable the selection of interpretable and traceable negative samples which can dynamically represent the most suitable instances for recommender. Two kinds of causality widely exist in RS, user-aspect, and interaction-aspect (Gao et al., 2022b). Regarding the user-aspect causality, user-similarity DNS can assist in unearthing commonalities within user behaviors by incorporating causal relationships, thus simulating the driving process behind user decision-making (Wu et al., 2019b; Wang et al., 2021b; Chen et al., 2019b). In contrast, concerning interaction-aspect causality, universal DNS and distribution-based DNS are proficient at reinforcing users’ unadulterated preferences by exploring and establishing causal links between user behaviors and recommended results (Shi et al., 2023; Zhu et al., 2022; Ding et al., 2020). Under these circumstances, negative sample selection is not solely based on the item correlation but also involves the discernment of potential causal relationships among them (Chen et al., 2019c). Nevertheless, the integration of negative sampling with causal inference remains an underexplored research avenue in RS, with only a limited number of studies delving into this domain. We believe that it holds great potential to offer the opportunity to provide more accurate, interpretable, and user-centric recommendations.
@@ 5.4. Alleviating Biases in Negative Sampling
To address this issue, some studies attempt to explore debiased learning by adversarial training (Krishnan et al., 2018; Anelli et al., 2021), causal modeling (Zheng et al., 2020; Wu et al., 2019e) and negative sampling (Ding et al., 2019; Chen et al., 2023b). It is able to improve the fairness of RS by rectifying the impact of biases, whether they stem from data imbalances, user historical behaviors, or various other factors (Chen et al., 2023a). Existing negative sampling algorithms primarily concentrate on mitigating exposure bias, popularity bias, and unfairness, which is achieved by over-sampling popular items (Gantner et al., 2012; Cheng et al., 2021; He et al., 2016) or enhancing the sampler with side information to facilitate model debiasing (Chen et al., 2023b; Steck, 2018; Zhao et al., 2019). Nevertheless, they have not proposed solutions for the broader issues of selection bias, conformity bias, and other biases in RS. The adaptive negative sampling strategies ensure that the learning process accounts for data imbalances and mitigates any potential biases in RS. Investigating how various negative sampling strategies affect the inherent biases in recommenders would be a genuinely interesting and valuable endeavor.
@@ 5.5. Incorporating Large Language Model into Recommendation
The rapid developments of Large Language Model (LLM) in Natural Language Processing have recently gained significant attention and revolutionized the learning paradigm of diverse research fields (Huang and Chang, 2022; Zhao et al., 2023). Some studies have indicated that the model size of LLM continues to expand with the immense scale of the corpus size, concomitantly enhancing their capabilities in logical reasoning and generalization (Hadi et al., 2023). Benefits from the potential to bridge the gap between the down-streaming task and the open-world knowledge, LLM have been incorporated into recommendation to serve as the knowledge extractor (Wu et al., 2021b; Hou et al., 2023a, 2022) and the recommender (Dai et al., 2023; Gao et al., 2023b; Hou et al., 2023b; Kang et al., 2023; Liu et al., 2023). Most of the existing LLM-based recommendation algorithms typically employ random sampling (Bao et al., 2023; hao et al., 2023; Mysore et al., 2023; Geng et al., 2022; Du et al., 2024; Wu et al., 2024) or leverage users’ explicit scores (Lin et al., 2024; Xi et al., 2023; Dai et al., 2023; Kang et al., 2023; Li et al., 2023; Lin et al., 2024) in real-world RS (Ratings within the range of are regarded as positive feedback, while defining the rest as negative feedback) for training. The negative samples selected in M6Rec (Cui et al., 2022) are exclusively from real-world systems, that is, Taobao and Alipay. To prevent the excessive training time and the overwhelming influx of negative information that could impact the model’s training process, M6Rec further down-samples the negative samples by a factor of ten. PALR (Yang et al., 2023a) incorporates the negative samples which are similar to the positive samples (e.g., movies belonging to the same genres and have been co-watched by many users) into model optimization.
@@
From the comprehensive review of existing LLM-based recommendation studies, we notice that LLMRank (Hou et al., 2023b) first investigates the ranking performance of LLM on hard negative candidates, which are retrieved by different classical recommendation algorithms. Moreover, InstructRec (Zhang et al., 2023) has also delved into LLM’s effectiveness and robustness in tasks involving more challenging hard negative samples and larger candidate sets. Nevertheless, these LLM-based recommendation algorithms have not taken into account negative sampling. Furthermore, the prevalent paradigm in these algorithms lies in employing natural language for depicting user interests, context, task formats, and historical behaviors. However, the omission of users’ negative preferences in this paradigm introduces the potential bias during the training and inference of LLM, thereby impacting its downstream recommendation performance. A lot of research has demonstrated that the adoption of well-reasoned negative sampling techniques in RS can further bolster the recommenders’ performance (Shi et al., 2023; Zhou et al., 2021; Lian et al., 2020a; Wu et al., 2021a). Negative Sampling would serve as a bridge between LLM and recommendation algorithms, facilitating the smooth incorporation of LLM into recommendation tasks. This contributes to being a powerful tool for assisting future researchers in constructing a LLM-based general recommendation framework.
@@ Acknowledgements.
This work is supported in part by the TaiShan Scholars Program (Grant no. tsqn202211289), the Shandong Province Excellent Young Scientists Fund Program (Overseas) (Grant no. 2022HWYQ-048), the Oversea Innovation Team Project of the ”20 Regulations for New Universities” funding program of Jinan (Grant no. 2021GXRC073) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). ChatGPT and Grammarly were utilized to improve grammar and correct spelling.References
@@
- Bao et al. (2023) Keqin Bao, Jizhi Zhang, Wenjie Wang, Yang Zhang, Zhengyi Yang, Yancheng Luo, Fuli Feng, Xiangnaan He, and Qi Tian. 2023. A Bi-Step Grounding Paradigm for Large Language Models in Recommendation Systems. (2023). http://arxiv.org/abs/2308.08434
@@
- Cai et al. (2022) Borui Cai, Yong Xiang, Longxiang Gao, He Zhang, Yunfeng Li, and Jianxin Li. 2022. Temporal knowledge graph completion: A survey. arXiv preprint arXiv:2201.08236 (2022).
@@
- Chen et al. (2023b) Xiao Chen, Wenqi Fan, Jingfan Chen, Haochen Liu, Zitao Liu, Zhaoxiang Zhang, and Qing Li. 2023b. Fairly Adaptive Negative Sampling for Recommendations. arXiv preprint arXiv:2302.08266 (2023).
@@
- Dai and Lin (2017) Bo Dai and Dahua Lin. 2017. Contrastive learning for image captioning. Advances in Neural Information Processing Systems 30 (2017).
@@
- Fan et al. (2023) Wenqi Fan, Zihuai Zhao, Jiatong Li, Yunqing Liu, Xiaowei Mei, Yiqi Wang, Zhen Wen, Fei Wang, Xiangyu Zhao, Jiliang Tang, and Qing Li. 2023. Recommender Systems in the Era of Large Language Models (LLMs). (2023). http://arxiv.org/abs/2307.02046
@@
- Gao et al. (2023a) Chongming Gao, Kexin Huang, Jiawei Chen, Yuan Zhang, Biao Li, Peng Jiang, Shiqi Wang, Zhong Zhang, and Xiangnan He. 2023a. Alleviating matthew effect of offline reinforcement learning in interactive recommendation. arXiv preprint arXiv:2307.04571 (2023).
- Gao et al. (2021a) Chen Gao, Yong Li, Fuli Feng, Xiangning Chen, Kai Zhao, Xiangnan He, and Depeng Jin. 2021a. Cross-domain recommendation with bridge-item embeddings. ACM Transactions on Knowledge Discovery from Data (TKDD) 16, 1 (2021), 1–23.
@@
- Gao et al. (2022b) Chen Gao, Yu Zheng, Wenjie Wang, Fuli Feng, Xiangnan He, and Yong Li. 2022b. Causal inference in recommender systems: A survey and future directions. arXiv preprint arXiv:2208.12397 (2022).
@@
- Gao et al. (2021b) Tianyu Gao, Xingcheng Yao, and Danqi Chen. 2021b. Simcse: Simple contrastive learning of sentence embeddings. arXiv preprint arXiv:2104.08821 (2021).
- Gao et al. (2022a) Yunjun Gao, Yuntao Du, Yujia Hu, Lu Chen, Xinjun Zhu, Ziquan Fang, and Baihua Zheng. 2022a. Self-guided learning to denoise for robust recommendation. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 1412–1422.
@@
- Giobergia (2022) Flavio Giobergia. 2022. Triplet Losses-based Matrix Factorization for Robust Recommendations. arXiv preprint arXiv:2210.12098 (2022).
@@
- Hassani and Khasahmadi (2020) Kaveh Hassani and Amir Hosein Khasahmadi. 2020. Contrastive multi-view representation learning on graphs. In International conference on machine learning. PMLR, 4116–4126.
@@
- Hou et al. (2023a) Yupeng Hou, Zhankui He, Julian McAuley, and Wayne Xin Zhao. 2023a. Learning vector-quantized item representation for transferable sequential recommenders. In Proceedings of the ACM Web Conference 2023.
@@
- Hou et al. (2023b) Yupeng Hou, Junjie Zhang, Zihan Lin, Hongyu Lu, Ruobing Xie, Julian McAuley, and Wayne Xin Zhao. 2023b. Large Language Models are Zero-Shot Rankers for Recommender Systems. (2023). http://arxiv.org/abs/2305.08845
@@
- Huang and Chang (2022) Jie Huang and Kevin Chen-Chuan Chang. 2022. Towards reasoning in large language models: A survey. arXiv preprint arXiv:2212.10403 (2022).
@@
- Kim et al. (2021b) Taeuk Kim, Kang Min Yoo, and Sang-goo Lee. 2021b. Self-guided contrastive learning for BERT sentence representations. arXiv preprint arXiv:2106.07345 (2021).
@@
- Lai et al. (2024) Riwei Lai, Rui Chen, Qilong Han, Chi Zhang, and Li Chen. 2024. Adaptive hardness negative sampling for collaborative filtering. arXiv preprint arXiv:2401.05191 (2024).
@@
- Li et al. (2023) Xiangyang Li, Bo Chen, Lu Hou, and Ruiming Tang. 2023. CTRL: Connect Collaborative and Language Model for CTR Prediction. (2023). http://arxiv.org/abs/2306.02841

### v2

@@ 4.5. Cross-domain Recommendation
Cross-domain recommendation (CDR) is one of the representative methods to alleviate the data sparsity problem and the cold-start problems in traditional RS with the auxiliary user behaviors collected from other domains (Cao et al., 2022; Ma et al., 2024b). It aims to leverage the correlation of users’ multi-domain behaviors to accurately transfer the informative knowledge from the source domains and model users’ personalized preference in the target domain (Bi et al., 2020). Traditional CDR methods typically conduct the cross-domain knowledge transfer by modeling the feature-level cross-domain correlations with aligned constraint (Krishnan et al., 2020; Liu et al., 2020), adversarial training (Liu et al., 2021a) and contrastive learning (Zheng et al., 2022; Ma et al., 2024b). However, these methods merely focus on the randomly-selected negative samples from the target domain, ignoring the cross-domain differences at the sample level. This may disregard users’ preference discrepancy between multiple domains and leads to the sub-optimal performance. DDGHM (Zheng et al., 2022) utilizes random mask operator for negative pair augmentation in CL optimization. Nevertheless, when it comes to recommendation tasks, the approach of random sampling within the target domain is still employed. RealHNS (Ma et al., 2023) is the pioneer to systematically discover the false and refine the real from all HNS in CDR. It conducts the multi-grained HNS selectors to find high-quality general and cross-domain HNS with a curriculum learning framework to alleviate cross-domain negative transfer. Notably, there exists substantial uncharted territories for negative sampling methods in CDR, which awaits the investigation by future researchers. This encompasses various areas, including, but not limited to, KNS, ANG, and Debiased IRW.
@@ 4.6. CL-enhanced Recommendation
Inspired by the success of CL in computer vision (Han et al., 2021), natural language processing (Kim et al., 2021b) and graph representation learning (Mo et al., 2022), some works (Shuai et al., 2022; Wang et al., 2022; Zhou et al., 2020; Tang et al., 2024a) apply CL paradigm in RS. CL aims to maximize the agreement between positive pairs (an item’s representation and its corresponding augmented representations) and minimize it between negative pairs (an item’s representation and other items’ augmented representations) to acquire the transferable knowledge between the original view and the augmented view. Intuitively, CL relies heavily on its negative augmentation (NA) strategy, which is due to the fact that the appropriate augmented view can facilitate the recommender to exploit richer underlying semantic correlation. The most widely used augmentation strategy in CL-based recommendation is the in-batch negative sampling. It treats the augmented representations of the remaining items in the mini-batch where the positive example is located as negative examples, rather than selecting from the corpus, thereby aiding in accelerating its training process without the imposition of additional computational complexity.
@@
To deeply understand the negative augmentation strategies in CL-based recommendation, we comprehensively dissect its core methodology into four distinct dimensions: pre-defined NA, graph-based NA, sequence-based NA and scenario-aware NA. Firstly, pre-defined NA fundamentally aims to refine CL process by augmenting the negative sample pool based on the pre-defined criteria (Du et al., 2022; Zhang et al., 2021b; Gu et al., 2022), which are typically derived from task-specific expertise. For instance, CauseRec (Zhang et al., 2021b) models the counterfactual data distribution by replacing the identified indispensable concepts in user behavioral sequences as negative pairs. Secondly, graph-based NA refers to generating the negative pairs based on distinct types of graph structures, including interaction graphs, social graphs, and heterogeneous graphs (Zhang et al., 2021a; Wu et al., 2021a; Yang et al., 2021d; Ma et al., 2022). It typically perturbs the nodes and edges to provide a nuanced understanding of user-item interactions. SGL (Wu et al., 2021a), EGLN (Yang et al., 2021d), and RGCL (Shuai et al., 2022) fully consider the correlations between nodes and edges within the graph structure to propose several graph-based NA strategies, including node (edge) dropout, node (edge) discrimination, random walk, and structure perturbation. Thirdly, sequence-based NA focuses on augmenting negative samples (e.g., shuffle, mask, crop, etc.) derived from historical behavior sequences (Zhou et al., 2020; Tian et al., 2022; Xie et al., 2022a; Wang et al., 2022). Its effective mechanism resides in comprehensively capturing the inherent dynamic characteristics of user behavior sequences. With the tailored self-supervised representation augmentation, it effectively handles the sequence structure and builds more robust and accurate recommenders. Finally, scenario-aware NA involves tailoring the augmentation of negative samples to accommodate the intricacies and demands of various recommendation scenarios. Here, MMR derives multi-view relationships of the user’s preferences and item’s representation across multiple modalities with CL to avoid the problem of limited diversity or bias (Wei et al., 2021; Tao et al., 2022), and MBR constructs negative examples based on the auxiliary behavior data to model the different user behavior patterns through CL (Gu et al., 2021; Yang et al., 2021a). Adaptively selecting the appropriate NA strategy based on the scenario requirements and data availability in CL-based recommendation is sensible. Meanwhile, how to construct a universal augmentation method to serve various recommendation scenarios remains an urgent issue.
@@ 5.3. Causal Inference for Negative Sample Understanding
Causality is the science of cause and effect in which the effect partly originates from the cause to some extent. As mentioned in (Gao et al., 2024), causal inference refers to the process of determining and further leveraging the causal relation from the experimental results or observational data. In the recent past, causal inference has been introduced into RS (Yang et al., 2021b) to extract the causality within the recommenders in modeling intricate user preferences, departing from the traditional correlation-centric paradigm (He et al., 2022, 2023b). This paradigm aims to determine how user historical behaviors influence the recommendation results, allowing for a comprehensive comprehension of user preferences.
From the perspective of negative sampling, we posit that the integration of negative sampling with causal inference holds the potential to enable the selection of interpretable and traceable negative samples which can dynamically represent the most suitable instances for recommender. Two kinds of causality widely exist in RS, user-aspect, and interaction-aspect (Gao et al., 2024). Regarding the user-aspect causality, user-similarity DNS can assist in unearthing commonalities within user behaviors by incorporating causal relationships, thus simulating the driving process behind user decision-making (Wu et al., 2019b; Wang et al., 2021b; Chen et al., 2019b). In contrast, concerning interaction-aspect causality, universal DNS and distribution-based DNS are proficient at reinforcing users’ unadulterated preferences by exploring and establishing causal links between user behaviors and recommended results (Shi et al., 2023; Zhu et al., 2022; Ding et al., 2020). Under these circumstances, negative sample selection is not solely based on the item correlation but also involves the discernment of potential causal relationships among them (Chen et al., 2019c). Nevertheless, the integration of negative sampling with causal inference remains an underexplored research avenue in RS, with only a limited number of studies delving into this domain. We believe that it holds great potential to offer the opportunity to provide more accurate, interpretable, and user-centric recommendations.
@@ 5.4. Alleviating Biases in Negative Sampling
To address this issue, some studies attempt to explore debiased learning by adversarial training (Krishnan et al., 2018; Anelli et al., 2021), causal modeling (Zheng et al., 2021; Wu et al., 2019d) and negative sampling (Ding et al., 2019; Chen et al., 2023b). It is able to improve the fairness of RS by rectifying the impact of biases, whether they stem from data imbalances, user historical behaviors, or various other factors (Chen et al., 2023a). Existing negative sampling algorithms primarily concentrate on mitigating exposure bias, popularity bias, and unfairness, which is achieved by over-sampling popular items (Gantner et al., 2012; Cheng et al., 2021; He et al., 2016) or enhancing the sampler with side information to facilitate model debiasing (Chen et al., 2023b; Steck, 2018; Zhao et al., 2019). Nevertheless, they have not proposed solutions for the broader issues of selection bias, conformity bias, and other biases in RS. The adaptive negative sampling strategies ensure that the learning process accounts for data imbalances and mitigates any potential biases in RS. Investigating how various negative sampling strategies affect the inherent biases in recommenders would be a genuinely interesting and valuable endeavor.
@@ 5.5. Incorporating Large Language Model into Recommendation
The rapid developments of Large Language Model (LLM) in Natural Language Processing have recently gained significant attention and revolutionized the learning paradigm of diverse research fields (Huang and Chen-Chuan Chang, 2023; Zhao et al., 2023; Fang et al., 2024; Jiang et al., 2025). Some studies have indicated that the model size of LLM continues to expand with the immense scale of the corpus size, concomitantly enhancing their capabilities in logical reasoning and generalization (Hadi et al., 2023). Benefits from the potential to bridge the gap between the down-streaming task and the open-world knowledge, LLM have been incorporated into recommendation to serve as the knowledge extractor (Wu et al., 2021b; Hou et al., 2023, 2022) and the recommender (Dai et al., 2023; Gao et al., 2023b; Hou et al., 2024; Kang et al., 2023; Liu et al., 2023b; Ma et al., 2025). Most of the existing LLM-based recommendation algorithms typically employ random sampling (Bao et al., 2025; hao et al., 2023; Mysore et al., 2023; Geng et al., 2022; Du et al., 2024; Wu et al., 2024d) or leverage users’ explicit scores (Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Kang et al., 2023; Li et al., 2023; Lin et al., 2024) in real-world RS (Ratings within the range of are regarded as positive feedback, while defining the rest as negative feedback) for training. The negative samples selected in M6Rec (Cui et al., 2022) are exclusively from real-world systems, that is, Taobao and Alipay. To prevent the excessive training time and the overwhelming influx of negative information that could impact the model’s training process, M6Rec further down-samples the negative samples by a factor of ten. PALR (Yang et al., 2023a) incorporates the negative samples which are similar to the positive samples (e.g., movies belonging to the same genres and have been co-watched by many users) into model optimization.
@@
From the comprehensive review of existing LLM-based recommendation studies, we notice that LLMRank (Hou et al., 2024) first investigates the ranking performance of LLM on hard negative candidates, which are retrieved by different classical recommendation algorithms. Moreover, InstructRec (Zhang et al., 2023a) has also delved into LLM’s effectiveness and robustness in tasks involving more challenging hard negative samples and larger candidate sets. Nevertheless, these LLM-based recommendation algorithms have not taken into account negative sampling. Furthermore, the prevalent paradigm in these algorithms lies in employing natural language for depicting user interests, context, task formats, and historical behaviors. However, the omission of users’ negative preferences in this paradigm introduces the potential bias during the training and inference of LLM, thereby impacting its downstream recommendation performance. A lot of research has demonstrated that the adoption of well-reasoned negative sampling techniques in RS can further bolster the recommenders’ performance (Shi et al., 2023; Zhou et al., 2021; Lian et al., 2020a; Wu et al., 2021a). Negative Sampling would serve as a bridge between LLM and recommendation algorithms, facilitating the smooth incorporation of LLM into recommendation tasks. This contributes to being a powerful tool for assisting future researchers in constructing a LLM-based general recommendation framework.
@@ Acknowledgements.
This work is supported in part by the TaiShan Scholars Program (Grant no. tsqn202211289), the Shandong Province Excellent Young Scientists Fund Program (Overseas) (Grant no. 2022HWYQ-048), the Oversea Innovation Team Project of the ”20 Regulations for New Universities” funding program of Jinan (Grant no. 2021GXRC073) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).References
@@
- Bao et al. (2025) Keqin Bao, Jizhi Zhang, Wenjie Wang, Yang Zhang, Zhengyi Yang, Yanchen Luo, Chong Chen, Fuli Feng, and Qi Tian. 2025. A Bi-Step Grounding Paradigm for Large Language Models in Recommendation Systems. ACM Trans. Recomm. Syst. 3, 4 (2025), 27 pages. https://doi.org/10.1145/3716393
@@
- Cai et al. (2023) Borui Cai, Yong Xiang, Longxiang Gao, He Zhang, Yunfeng Li, and Jianxin Li. 2023. Temporal knowledge graph completion: a survey. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence (IJCAI ’23). https://doi.org/10.24963/ijcai.2023/734
@@
- Chen et al. (2023e) Jiawei Chen, Junkang Wu, Jiancan Wu, Xuezhi Cao, Sheng Zhou, and Xiangnan He. 2023e. Adap-: Adaptively Modulating Embedding Magnitude for Recommendation. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). 1085–1096. https://doi.org/10.1145/3543507.3583363
@@
- Chen et al. (2023b) Xiao Chen, Wenqi Fan, Jingfan Chen, Haochen Liu, Zitao Liu, Zhaoxiang Zhang, and Qing Li. 2023b. Fairly Adaptive Negative Sampling for Recommendations. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). 3723–3733. https://doi.org/10.1145/3543507.3583355
@@
- Chen et al. (2024) Yuxin Chen, Junfei Tan, An Zhang, Zhengyi Yang, Leheng Sheng, Enzhi Zhang, Xiang Wang, and Tat-Seng Chua. 2024. On Softmax Direct Preference Optimization for Recommendation. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 27463–27489. https://proceedings.neurips.cc/paper_files/paper/2024/file/30732ddb12d9faf7180f5d0e8b5b5da7-Paper-Conference.pdf
@@
- Dai et al. (2024a) Sunhao Dai, Changle Qu, Sirui Chen, Xiao Zhang, and Jun Xu. 2024a. ReCODE: Modeling Repeat Consumption with Neural ODE. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval. 2599–2603.
@@
- Dai et al. (2024b) Sunhao Dai, Ninglu Shao, Jieming Zhu, Xiao Zhang, Zhenhua Dong, Jun Xu, Quanyu Dai, and Ji-Rong Wen. 2024b. Modeling User Attention in Music Recommendation. In 2024 IEEE 40th International Conference on Data Engineering (ICDE). IEEE, 761–774.
@@
- Fan (2024) Wenqi Fan. 2024. Recommender systems in the era of large language models (llms). IEEE Transactions on Knowledge and Data Engineering (2024), 1–20.
- Fang et al. (2024) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-seng Chua. 2024. Alphaedit: Null-space constrained knowledge editing for language models. arXiv preprint arXiv:2410.02355 (2024).
@@
- Gao et al. (2023a) Chongming Gao, Kexin Huang, Jiawei Chen, Yuan Zhang, Biao Li, Peng Jiang, Shiqi Wang, Zhong Zhang, and Xiangnan He. 2023a. Alleviating matthew effect of offline reinforcement learning in interactive recommendation. In Proceedings of the 46th international ACM SIGIR conference on research and development in information retrieval. 238–248.
@@
- Gao et al. (2024) Chen Gao, Yu Zheng, Wenjie Wang, Fuli Feng, Xiangnan He, and Yong Li. 2024. Causal inference in recommender systems: A survey and future directions. ACM Transactions on Information Systems 42, 4 (2024), 1–32.
@@
- Gao et al. (2022) Yunjun Gao, Yuntao Du, Yujia Hu, Lu Chen, Xinjun Zhu, Ziquan Fang, and Baihua Zheng. 2022. Self-guided learning to denoise for robust recommendation. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 1412–1422.
@@
- Giobergia (2022) Flavio Giobergia. 2022. Triplet Losses-based Matrix Factorization for Robust Recommendations. CoRR abs/2210.12098 (2022). https://doi.org/10.48550/ARXIV.2210.12098 arXiv:2210.12098
@@
- Ha et al. (2024) Mingming Ha, Xuewen Tao, Wenfang Lin, Qiongxu Ma, Wujiang Xu, and Linxun Chen. 2024. Fine-Grained Dynamic Framework for Bias-Variance Joint Optimization on Data Missing Not at Random. Advances in Neural Information Processing Systems 37 (2024), 104010–104034.
@@
- Hou et al. (2023) Yupeng Hou, Zhankui He, Julian McAuley, and Wayne Xin Zhao. 2023. Learning vector-quantized item representation for transferable sequential recommenders. In Proceedings of the ACM Web Conference 2023.
@@
- Hou et al. (2024) Yupeng Hou, Junjie Zhang, Zihan Lin, Hongyu Lu, Ruobing Xie, Julian McAuley, and Wayne Xin Zhao. 2024. Large language models are zero-shot rankers for recommender systems. In European Conference on Information Retrieval. Springer, 364–381.
@@
- Huang and Chen-Chuan Chang (2023) Jie Huang and Kevin Chen-Chuan Chang. 2023. Towards Reasoning in Large Language Models: A Survey. In Findings of the Association for Computational Linguistics, ACL 2023 (Proceedings of the Annual Meeting of the Association for Computational Linguistics). Association for Computational Linguistics (ACL), 1049–1065. https://doi.org/10.18653/v1/2023.findings-acl.67
@@
- Jiang et al. (2025) Houcheng Jiang, Junfeng Fang, Ningyu Zhang, Guojun Ma, Mingyang Wan, Xiang Wang, Xiangnan He, and Tat-seng Chua. 2025. AnyEdit: Edit Any Knowledge Encoded in Language Models. arXiv preprint arXiv:2502.05628 (2025).
@@
- Kim et al. (2021b) Taeuk Kim, Kang Min Yoo, and Sang goo Lee. 2021b. Self-Guided Contrastive Learning for BERT Sentence Representations. ArXiv abs/2106.07345 (2021). https://api.semanticscholar.org/CorpusID:235422673
@@
- Lai et al. (2024) Riwei Lai, Rui Chen, Qilong Han, Chi Zhang, and Li Chen. 2024. Adaptive hardness negative sampling for collaborative filtering. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 8645–8652.
@@
- Li et al. (2024) Haoxuan Li, Chunyuan Zheng, Wenjie Wang, Hao Wang, Fuli Feng, and Xiao-Hua Zhou. 2024. Debiased recommendation with noisy feedback. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 1576–1586.
@@
- Li et al. (2023) Xiangyang Li, Bo Chen, Lu Hou, and Ruiming Tang. 2023. Ctrl: Connect collaborative and language model for ctr prediction. ACM Transactions on Recommender Systems (2023).